# Where did heavy binaries go? Gravitational-wave populations using Delaunay triangulation with optimized complexity

### v2

@@ 4 Conclusions
So, where would the heavy binaries have gone? Taken together, our results favor a mixed population at $M_{\mathrm{tot}}\sim 70\,\mathrm{M}_{\odot}$ with at least two components: a short–time-delay channel stemming from dense environments that dominates at high redshift, and a longer–time-delay component consistent with isolated (and ejected) binaries that becomes dominant toward low redshift.
Moreover, our findings imply that analyses measuring cosmological parameters from the redshift dependence of the detector-frame mass spectrum (Chernoff & Finn, 1993; Markovic, 1993; Farr et al., 2019; Mastrogiovanni et al., 2021; Abac et al., 2025b) should carefully account for the evolution of the prominence of the high-mass peak with $z$.
The key strength of using Delaunay triangulation is that the interpolation skeleton, including the number and location of the nodes, is inferred directly from the data. In practice, this optimized complexity reduces the number of parameters by one to two orders of magnitude compared to the pixelized approach by Heinzel et al. (2025b), and attain a good recovery of arbitrary functional forms at a lower tuning cost compared to the binned Gaussian-process approach by Ray et al. (2023), both of which rely on a fixed grid. For this reason, we expect our flexible framework to be significantly more efficient at probing correlations in dimensions $\geq 3$, making it ideally suited for the forthcoming big-data era of GW astronomy.
@@ Appendix A Delaunay triangulation for inference
We seek to represent general multidimensional distributions using an agnostic model with a low computational cost. Our strategy is to construct a trans-dimensional interpolator for the merger rate using barycentric interpolation on a Delaunay triangulation (Delaunay, 1934; Rippa, 1990; Rajan, 1994). The number of triangulation points is allowed to vary through inference, thus modeling arbitrarily simple (or complex) distributions depending on the available information in the data. This method can be applied to parameter spaces of arbitrary dimensionality. While we restrict our application of the method to a two-dimensional parameter space, we describe it in full generality.
@@
We denote the (natural log) differential rate as computed by a triangulation with parameters $\Lambda$ by $\log\mathrm{d}_{\theta}N(\theta|\Lambda)\equiv\Delta(\theta|\Lambda)$. To compute the differential rate at a given point $\theta$, we find the simplex $S(\theta)$ (i.e. triangle if $D=2$) wherein it lies, compute its barycentric coordinates, and use them to linearly interpolate the weights at their vertices.
@@
We construct Delaunay triangulations using qhull (Barber et al., 1996) via scipy (Virtanen et al., 2020).
@@ Appendix B Mass-ratio distribution
The fiducial model by Abac et al. (2025c) describes the joint $(m_{1},q)$ distribution as a sum of components, each independent in $m_{1}$ and $z$. This formulation makes the overall distribution non-separable in $m_{1}$ and $q$, and this cannot be straightforwardly implemented in our framework. Guided by Abac et al. (2025c)’s findings (see their Fig. 5), we describe the mass-ratio distribution using a broken power law:
@@ Appendix C Formalism for hierarchical inference
We express the population likelihood for observing $N_{\mathrm{obs}}$ events with data $\{d_{j}\}=\{d_{1},\dots,d_{N_{\mathrm{obs}}}\}$ for a set of population hyperparameters $\Lambda$ as (Mandel et al., 2019; Vitale et al., 2020)
@@
For cosmological calculations, we assume the Planck15 cosmology as implemented in astropy (Robitaille et al., 2013).
@@
Equations (C1) and (C2) are computed via Monte Carlo integration, here indicated with $\langle\cdot\rangle$. Selection effects are estimated using the injection campaign presented by Essick et al. (2025). We set the population likelihood to zero for hyperparameters that yield a variance in the Monte Carlo estimators greater than one (Talbot & Golomb, 2023; Abac et al., 2025c; Mancarella & Gerosa, 2025). We verify the robustness of this choice by re-running our inference using thresholds of 0.8 and 1.2. The results are qualitatively unchanged; the credibility for an enhanced merger rate at $\sim 70\,\mathrm{M}_{\odot}$ is $\sim 95\%$ instead of $\sim 98\%$ for the 1.2 threshold.
@@
As discussed in the main text, the prior on the modeled part of the population parameters $(\Lambda_{q},\Lambda_{\chi},\Lambda_{\vartheta})$ are taken equal to those of Abac et al. (2025c), while the prior on the triangulation parameters is uniform across the area of interest for the triangulation vertices and along (20,15)(-20,15) for the triangulation weights. The implied prior range on $\mathrm{d}_{\theta}N$ (or equivalently $\mathrm{d}_{\theta}\mathcal{R}$) is compatible with (if not more conservative than) those used in similar non-parametric approaches (Heinzel et al., 2025b, a; Alvarez-Lopez et al., 2025). We also verified the robustness of our results against prior choices by extending the weight prior down to (25,15)(-25,15), which lowers the induced prior on $\mathrm{d}_{\theta}N$ by two to three orders of magnitude. The results for $z=0.2$ remain unchanged. For $z=1$, the lower end of the posterior distribution on $\mathrm{d}_{\theta}\mathcal{R}$ lowers so that the credibility for $\mathrm{d}_{\theta}\mathcal{R}(z=1)>\mathrm{d}_{\theta}\mathcal{R}(z=0.2)$ is slightly lower ($\sim 94\%$ instead of $\sim 98\%$). We stress that the key result of this work, namely the disappearance of the peak at low redshifts, is not affected by prior choices. The lack of an obvious choice for the prior range of $w_{i}$ suggests non-parametric population models are better suited to conduct inference in the observed space, where the domain of inference is naturally limited to data-dominated regions (Toubiana et al., 2025).
@@
We also compute the Kullback-Leibler divergence $D_{\mathrm{KL}}$ (Kullback & Leibler, 1951) to compare the posterior volumetric merger rate to its prior distribution at any given $(M_{\mathrm{tot}},z)$. The prior distribution on $\log_{10}\mathrm{d}_{\theta}\mathcal{R}(M_{\mathrm{tot}},z)$, denoted by $p(\log_{10}\mathrm{d}_{\theta}\mathcal{R})$ for simplicity, is constructed by sampling the prior distribution of $\Lambda_{\Delta}$ (i.e. number of vertices, positions, and weights) and computing the corresponding $\log_{10}\mathrm{d}_{\theta}\mathcal{R}(M_{\mathrm{tot}},z)$. The posterior distribution, denoted $p(\log_{10}\mathrm{d}_{\theta}\mathcal{R}|\{d_{j}\})$, is computed analogously by sampling from $p(\Lambda_{\Delta}|\{d_{j}\})$. To approximate these probability densities, we generate $10^{4}$ samples in every $(M_{\mathrm{tot}},z)$ of interest and fit a Gaussian Kernel Density Estimator. The Kullback-Leibler divergence (in nat) for any given $(M_{\mathrm{tot}},z)$ is then computed as

### v3

@@ 4 Conclusions
So, where would the heavy binaries have gone? Taken together, our results suggest the presence of a population with $M_{\mathrm{tot}}\sim 70\mathrm{M}_{\odot}$, characterized by a fast-merging component that dominates at high redshift and a slow-merging component that becomes increasingly important at low redshift. While such a trend could arise from multiple formation pathways—such as dense environments for short delays and isolated (and ejected) binaries for long delays—establishing the relative contribution of specific channels requires further observational and theoretical work. This includes investigating potential correlations in other parameters, most notably spins and eccentricity.
Moreover, our findings imply that analyses measuring cosmological parameters from the redshift dependence of the detector-frame mass spectrum (Chernoff:1993th; Markovic:1993cr; Farr:2019twy; Mastrogiovanni:2021wsd; LIGOScientific:2025jau) should carefully account for the evolution of the prominence of the high-mass peak with $z$.
The key strength of using Delaunay triangulation is that the interpolation skeleton, including the number and location of the nodes, is inferred directly from the data. In practice, this optimized complexity reduces the number of parameters by one to two orders of magnitude compared to the pixelized approach by Heinzel:2024hva, and attain a good recovery of arbitrary functional forms at a lower tuning cost compared to the binned Gaussian-process approach by Ray:2023upk, both of which rely on a fixed grid. For this reason, we expect our flexible framework to be significantly more efficient at probing correlations in dimensions $\geq 3$, making it ideally suited for the forthcoming big-data era of GW astronomy.
@@ Appendix A Delaunay triangulation for inference
We seek to represent general multidimensional distributions using an agnostic model with a low computational cost. Our strategy is to construct a trans-dimensional interpolator for the merger rate using barycentric interpolation on a Delaunay triangulation (Delaunay_1934aa; RIPPA1990489; rajan1994). The number of triangulation points is allowed to vary through inference, thus modeling arbitrarily simple (or complex) distributions depending on the available information in the data. This method can be applied to parameter spaces of arbitrary dimensionality. While we restrict our application of the method to a two-dimensional parameter space, we describe it in full generality.
@@
We denote the (natural log) differential rate as computed by a triangulation with parameters $\Lambda$ by $\log\mathrm{d}_{\theta}N(\theta|\Lambda)\equiv\Delta(\theta|\Lambda)$. To compute the differential rate at a given point $\theta$, we find the simplex $S(\theta)$ (i.e. triangle if $D=2$) wherein it lies, compute its barycentric coordinates, and use them to linearly interpolate the weights at their vertices:
@@
We construct Delaunay triangulations using qhull (Barber:1996lmi) via scipy (Virtanen:2019joe).
@@ Appendix B Mass-ratio distribution
The fiducial model by LIGOScientific:2025pvj describes the joint $(m_{1},q)$ distribution as a sum of components, each independent in $m_{1}$ and $z$. This formulation makes the overall distribution non-separable in $m_{1}$ and $q$, and this cannot be straightforwardly implemented in our framework. Guided by LIGOScientific:2025pvj’s findings (see their Fig. 5), we describe the mass-ratio distribution using a broken power law:
@@ Appendix C Formalism for hierarchical inference
We express the population likelihood for observing $N_{\mathrm{obs}}$ events with data $\{d_{j}\}=\{d_{1},\dots,d_{N_{\mathrm{obs}}}\}$ for a set of population hyperparameters $\Lambda$ as (Mandel:2018mve; Vitale:2020aaz)
@@
For cosmological calculations, we assume the Planck15 cosmology as implemented in astropy (Astropy:2013muo).
@@
Equations (C1) and (C2) are computed via Monte Carlo integration, here indicated with $\langle\cdot\rangle$. Selection effects are estimated using the injection campaign presented by Essick:2025zed. We set the population likelihood to zero for hyperparameters that yield a variance in the Monte Carlo estimators greater than one (Talbot:2023pex; LIGOScientific:2025pvj; Mancarella:2025uat). We verify the robustness of this choice by re-running our inference using thresholds of 0.8 and 1.2. The results are qualitatively unchanged; the credibility for an enhanced merger rate at $\sim 70\,\mathrm{M}_{\odot}$ is $\sim 95\%$ instead of $\sim 98\%$ for the 1.2 threshold.
@@
As discussed in the main text, the prior on the modeled part of the population parameters $(\Lambda_{q},\Lambda_{\chi},\Lambda_{\vartheta})$ are taken equal to those of LIGOScientific:2025pvj, while the prior on the triangulation parameters is uniform across the area of interest for the triangulation vertices and along (20,15)(-20,15) for the triangulation weights. The implied prior range on $\mathrm{d}_{\theta}N$ (or equivalently $\mathrm{d}_{\theta}\mathcal{R}$) is compatible with (if not more conservative than) those used in similar non-parametric approaches (Heinzel:2024hva; Heinzel:2024jlc; Alvarez-Lopez:2025ltt). We also verified the robustness of our results against prior choices by extending the weight prior down to (25,15)(-25,15), which lowers the induced prior on $\mathrm{d}_{\theta}N$ by two to three orders of magnitude. The results for $z=0.2$ remain unchanged. For $z=1$, the lower end of the posterior distribution on $\mathrm{d}_{\theta}\mathcal{R}$ lowers so that the credibility for $\mathrm{d}_{\theta}\mathcal{R}(z=1)>\mathrm{d}_{\theta}\mathcal{R}(z=0.2)$ is slightly lower ($\sim 94\%$ instead of $\sim 98\%$). We stress that the key result of this work, namely the disappearance of the peak at low redshifts, is not affected by prior choices. The lack of an obvious choice for the prior range of $w_{i}$ suggests non-parametric population models are better suited to conduct inference in the observed space, where the domain of inference is naturally limited to data-dominated regions (Toubiana:2025syw).
@@
We also compute the Kullback-Leibler divergence $D_{\mathrm{KL}}$ (kl_div) to compare the posterior volumetric merger rate to its prior distribution at any given $(M_{\mathrm{tot}},z)$. The prior distribution on $\log_{10}\mathrm{d}_{\theta}\mathcal{R}(M_{\mathrm{tot}},z)$, denoted by $p(\log_{10}\mathrm{d}_{\theta}\mathcal{R})$ for simplicity, is constructed by sampling the prior distribution of $\Lambda_{\Delta}$ (i.e. number of vertices, positions, and weights) and computing the corresponding $\log_{10}\mathrm{d}_{\theta}\mathcal{R}(M_{\mathrm{tot}},z)$. The posterior distribution, denoted $p(\log_{10}\mathrm{d}_{\theta}\mathcal{R}|\{d_{j}\})$, is computed analogously by sampling from $p(\Lambda_{\Delta}|\{d_{j}\})$. To approximate these probability densities, we generate $10^{4}$ samples in every $(M_{\mathrm{tot}},z)$ of interest and fit a Gaussian Kernel Density Estimator. The Kullback-Leibler divergence (in nat) for any given $(M_{\mathrm{tot}},z)$ is then computed as